Interim Report

Rhiza S. Sadjad*

Control Systems and Instrumentation Laboratory, Department of Electrical Engineering Faculty of Engineering, Hasanuddin University, MAKASSAR Indonesia 90245 Email: rhiza@unhas.ac.id, URL: http://www.unhas.ac.id/rhiza/

I. Introduction

This is an interim report (for the period of June 28 to September 28, 2012) on my participation in the research fellowship program related to the Hasanuddin University’s New Engineering Campus Development Project (JICA Loan No. IP-541) offered to interested faculty members and arranged in a 6 (six) month visit to a university in Japan. Thanks to the help of Prof. Manabu Kano (Kyoto University), Dr. Manabu Tsunoda (JICA) and the JICE, I was accepted to visit Prof. Hiroya Seki’s Process Systems Engineering Laboratory at Tokyo Institute of Technology (TITECH). I arrived at Tokyo on Thursday, June 28, 2012, settled in to stay at the TITECH’s International House at Ookayama TOKYO and met with Prof. Hiroya Seki in his laboratory at TITECH’s Suzekakedai campus YOKOHAMA, at the same day.

II. General Overview

Having 3 (three) different campuses: the Tamachi campus, the Ookayama campus and the Suzukakedai campus, TITECH (abreviated as ”Tokodai” in Japenese) could become a good model for the Hasanuddin University Faculty of Engineering, whose campus are also located in 3 (three) different places: the old Baraya campus, the current Tamalanrea campus and the new Gowa campus. The old Baraya campus could host the vocational, highschool level, extension and reach-out programs similar to the TITECH’s campus at Tamachi. The new Gowa campus could take the Ookayama campus as the model, emphasizing on the mono-disciplinary undergraduate and graduate programs. Inter-disciplinary and multi-disciplinary research and graduate programs could stay on in Tamalanrea campus to take the TITECH’s campus at Suzukakedai as the reference. By not moving those programs currently located at the Tamalanrea campus to Gowa campus, the Hasanuddin University Faculty of Engineering may still be able to take the advantage of being co-located with other Faculties in many inter-disciplinary and multi-disciplinary research and graduate programs. Geographically, the positions of the TITECH’s three campuses relative to each other are very similar to the positions of Hasanuddin University Faculty of Engineering’s three campuses. It takes around 40 minutes by train, for instance, to go from the Ookayama campus to the Suzukakedai campus, which is approximately the same length of time to go by car from the Tamalanrea campus to the new campus at Gowa.

III. Curriculum and Learning Process

I did not have a chance to take a look at the Tokodai’s curriculum and learning process in more details, yet from the limited talks and discusions I made with a few Indonesians who study at the Tokodai, I started to get an impression that the students at the Tokodai generally do not take many courses. Even the undergraduate students do not take many classroom courses, moreover the graduate students. As a matter of fact, I rarely see students attending lectures in the classrooms, especially in Tokodai’s campus at Suzukakedai. I saw the campuses are relatively very quiet, maybe because I arrived almost at the end of Spring semester and at the beginning of the summer vacation, while the next Fall semester will not begin until the end of September or the beginning of October. In many campuses in the world, especially

Fig. 1. The International at the TITECH’s Ookayama Campus

in Indonesia, we usually see university students are busy attending classes from time to time, from one building to another, but not here in Japan, at least not here in Tokodai.

The fact that in Tokodai the students spend more time in the laboratories rather than in the classrooms is very interesting. If the Hasanuddin University Faculty of Engineering is to follow this model, then its curriculum stucture should be changed from emphasizing on classroom courses like the current system into the new curriculum emphasizing on laboratory courses or research activities.

Fig. 2. The TITECH’s Suzukakedai Campus

In the future, when our department moves to the new campus approximately in 2013 or 2014, our laboratory will not only be able to support the academic program of our department - the Laboratory-Based Education (LBE) program - but more than that, it will take its role as the important part of an industrial training center serving the industrial community in the surrounding area, the eastern region of Indonesia. In order to enhance our services - in terms of academics, research and public services - we are now proposing to expand our department into a much wider scope of a faculty. The proposed name of the new organization is the Faculty of Electrical Engineering and Informatics. In the future, the new faculty is planned to establish at least 4 (four) study programs, i.e. the existing Electrical Engineering and Informatic Engineering, and two additional study programs, the Computer System Engineering and the Engineering Physics (specializing in the Process Control Technology and Instrumentation) study programs, both are to be expanded from the existing Computer, Control and Electronic Sub-study Program. With this future plan bare clearly in mind, for this 6-month visiting program in Japan, I have made an intention to focus on two issues, firstly, to observe the Japanese university-industry research collaboration, and secondly, to survey the current curriculum development and research activities of some prominent Japanese universities, especially both in the fields of process control technology and instrumentation systems. The main result of this visit is a draft of the proposal to prepare for the establishment of the Engineering Physics study program within the future Faculty of Electrical Engineering and Informatics.

Fig. 3. My Work Station at the Process Systems Engineering Laboratory

Japan has a long experience with the university-industry collaboration, especially in the field of process control technology [2] [3]. The short visit to both the university and the industrial world in Japan will be very beneficial for the future of our laboratory at its location in the new campus. The main concern is about the university-industry collaboration itself, another concern is related to the LBE curriculum development of the proposed Engineering Physics study program.

IV. Objectives

The main goal of the 6 (six) month visit to Japan is to collect the lessons learned from the Japanese experience in the collaboration between the academic world and the real industrial world, especially in the field of process control technology. The best model of university-industry collaboration will be developed as the main outcome of the visit, which is considered as the foundation of the establishment of the future Engineering Physics study program. The secondary goal is to survey the curriculum development and research activities of the Chemical Engineering and Engineering Physics Departments in some prominent universities in Japan, where the process control technology and instrumentation is their main competency. The visit will begin and end at the Tokyo Institute of Technology, Suzukakedai Campus at Yokohama, hosted by Professor Hiroya SEKI of the Process Systems Engineering Division, Chemical Resources Laboratory.

V. Schedule

Tentatively, the schedule of the visit can be arranged as shown in TABLE I. The first five days will be spent for orientation and preliminary discussions, fixing the details of the 6-month visiting program planning. For the next 2 months after my arrival I am planning to stay at the Suzukakedai Campus.

From the third to the fifth month I would like to make a tour to visit the surrounding industrial world, industrial research and training centers and universities, where the university-industry collaboration on the

TABLE I

The Tentative Schedule

Day 1 to Day 5  Orientation and preliminary discussion with the host professor at the TIT
June to July 2012  Staying and working at the TIT’s Chemical Resources Laboratory
August to October 2012 Visiting industrial plants, research centers, etc. in the surrounding area
November 2012  Back to the university, writing reports and papers

field of process control technology and instrumentation is in progress. The final sixth month will be spent for writing reports and papers to conclude the visit program.

References

[1] Andani Ahmad, The Miniature of an Industrial Solid Material Process Plant, Ph.D. Dissertation, Hasanuddin University Graduate Program, Makassar, Indonesia, 2010.

[2] Manabu Kano, Recent Development of Process Control Technology through Industry-University Collaboration in Japan , The 13th Asia Pacific Confederation of Chemical Engineering Congress , Taipei, Taiwan, October 5-8, 2010.

[3] Kano M. and Ogawa M., The State of the Art in Advanced Chemical Process Control in Japan, IFAC ADCHEM, CD-ROM, Istanbul, Turkey, July 12-15, 2009

* Rhiza S. Sadjad, was born in 1957, completed his elementary and secondary education at his home town Bogor, Indonesia. He received his first college degree of Ir. from the Department of Electrical Engineering, Bandung Institute of Technology (ITB), Bandung, Indonesia in 1981, then received the M.S.E.E. (1989) and the Ph.D. (1994) majoring in Automatic Control Systems from the Department of Electrical and Computer Engineering, University of Wisconsin, Madison WI, USA. In 1981 he took a teaching position at the Faculty of Electrical Engineering, Satya Wacana Christian University in Salatiga, Central Java, Indonesia, then moved to the Department of Electrical Engineering, Hasanuddin University, Makassar, South Sulawesi, Indonesia in 1983 and has been with this department until now.. He founded the Control Systems and Instrumentation Laboratory in 1995 and has been the head of the laboratory since then. He teaches almost all courses in automatic control systems area, and has recently been interested in the process control

systems and technology. He has supervised a Ph.D. dissertation (completed in 2010) to develop a process control mini-plant for industrial solid materials, and now is advising undergraduate and masters students whose final projects and thesis are related to the field of process control technology.